OpenStudy (anonymous):

State any restrictions on the variables. 3y^2-3y ------- y+1 A. y ≠ 1 B. y ≠ –1 C. y ≠ 3 D. no restrictions are needed

OpenStudy (anonymous):

@Perl @Mr.Expensive

OpenStudy (anonymous):

Though I cannot understand this problem very much, probably I will be like that. \[\frac{ 3^y{2}-3y }{ y+1 }=\frac{ 3(y^{2}-y) }{ y+1 }=\frac{ 3(y+1)(y-1) }{ y+1 }=3(y-1)\] \[\frac{ 3y^{2} -3y}{ y+1 }=0, y=1\] So the answer is "A".

OpenStudy (anonymous):

Thank you @Tsuyo

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Oh, don't worry about it.

OpenStudy (anonymous):

Actually, the answer was B...